OpenStudy (anonymous):

A geometric series has 3 terms. The sum of the 3 terms is 42. Term 3 is 3.2 times the sum of the other 2 terms. What are the terms?

OpenStudy (anonymous):

the first 2 terms are any number that add up to 10 and the last number is 32. you start with x+y+3.2(x+y)=42 distribut the 3.2 and combine like terms to get 4.2x+4.2y=42 then factor out the 4.2 : 4.2(x+y)=42 and divide by 4.2 x+y=10